### v1

@@ Abstract
This paper is concerned with the approximation of probability distributions known up to normalization constants, with a focus on Bayesian inference for large-scale inverse problems in scientific computing. In this context, key challenges include costly repeated evaluations of forward models, multimodality, and inaccessible gradients for the forward model. To address them, we develop a variational inference framework that combines Fisher-Rao natural gradient with specialized quadrature rules to enable derivative free updates of Gaussian mixture variational families. The resulting method, termed Derivative Free Gaussian Mixture Variational Inference (DF-GMVI), guarantees covariance positivity and affine invariance, offering a stable and efficient framework for approximating complex posterior distributions. The effectiveness of DF-GMVI is demonstrated through numerical experiments on challenging scenarios, including distributions with multiple modes, infinitely many modes, and curved modes in spaces with up to hundreds of dimensions. The method’s practicality is further demonstrated in a large-scale application, where it successfully recovers the initial conditions of the Navier-Stokes equations from solution data at positive times.
@@ 1 Introduction
Sampling a target probability distribution known up to normalization constants is a classical problem in scientific computing. Specifically, in Bayesian inverse problems [39, 57], the goal is to recover an unknown parameter $\theta\in\mathbb{R}^{N_{\theta}}$ from noisy observation $y\in\mathbb{R}^{N_{y}}$, through the equation
@@
Here, $\mathcal{G}$ represents a forward map which, for the problems we consider, is nonlinear and requires solving a partial differential equation (PDE) for each evaluation. The observational noise $\eta$ follows a Gaussian distribution: $\eta\sim\mathcal{N}(0,\Sigma_{\eta})$. Within the Bayesian framework, we assign a Gaussian prior $\mathcal{N}(r_{0},\Sigma_{0})$ to the unknown parameter $\theta$, resulting in a posterior distribution from which we aim to draw samples
@@
It is worth noting that in this context, $\Phi_{R}$ adopt a nonlinear least-squares structure, with an augmented map $\mathcal{F}(\theta)$ satisfying
@@
The primary goal of this paper is to develop an efficient approach for the approximate sampling of such $\rho_{\rm post}(\theta)$.
@@ 1.1 Challenges
For many Bayesian inverse problems in scientific applications, computing gradients of $\Phi_{R}$ might be infeasible, as it requires derivatives of large-scale PDE-based models $\mathcal{G}$ that may be black-box (e.g., climate models [55, 54]), use non-differentiable numerical methods (e.g., embedded boundary methods [51, 31, 30, 9] and adaptive mesh refinement [3, 6]), or model discontinuous physics (e.g., in fracture [48] or cloud modeling [58, 44]). Derivative free methods are thus needed; see a review of methodologies in Section 1.3.2. While there are a few Markov Chain Monte Carlo (MCMC) and Sequential Monte Carlo (SMC) approaches that do not require gradients, they often require numerous function evaluations, particularly in high dimensions, to achieve convergence or mitigate weight collapse. This makes them impractical given the high computational cost of evaluating $\mathcal{G}$. Furthermore, the multimodality of $\rho_{\rm post}(\theta)$ causes MCMC methods to struggle with mode transitions [23, 24]. We note that missing modes can lead to significant prediction errors in scientific applications [59].
Variational inference offers a promising alternative, with black-box variational inference (BBVI) [52] being a popular approach for achieving derivative free sampling. BBVI typically relies on stochastic approximations to estimate gradients which exist high variance. This often makes BBVI unreliable, requiring variance reduction techniques and extensive time step tuning. Even with these improvements, small time steps are necessary to maintain stability (see Section 5.2). Kalman methodology is also used to develop derivative-free variational inference methods [14, 32]. While these approaches can be effective for posterior distributions close to Gaussian or Gaussian mixture with separated components, they often become unstable in the presence of non-Gaussian or strongly curved modes (see [14, Appendix E]).
In this work, we explore guidelines for designing stable derivative free Gaussian mixture variational inference methods. Building on these guidelines, we propose a novel approach that requires no hyperparameter tuning while achieving both high accuracy and computational efficiency.
@@ 1.2 Contributions
    We propose derivative free quadrature rules for approximating the expectations of $\Phi_{R}$ and Gaussian mixtures, their gradients, and their Hessians with respect to any Gaussian density. These methods are exact for linear $\mathcal{F}$ implying low variance, and capture as much curvature information of $\mathcal{F}$ as possible. Moreover, the computational cost, in terms of evaluating the forward mapping $\mathcal{F}(\theta)$, scales linearly with the dimensionality of $\theta$.
@@
    Building on these quadrature rules, we develop a derivative free variational inference approach that combines Fisher-Rao natural gradient and Gaussian mixture approximations (DF-GMVI). The DF-GMVI method ensures covariance positivity and exhibits affine invariant properties, which together contribute to its superior stability even when we use large time steps. These guidelines can also be applied to improve the robustness of other Gaussian mixture variational inference methods.
@@
    We demonstrate that DF-GMVI effectively captures multiple or even infinite modes, as well as curved modes, in model problems involving up to hundreds of dimensions. Additionally, it performs well in PDE-based applications, such as reconstructing the Navier-Stokes initial condition from solution data at positive times.
@@ 1.3 Literature Review
The review of relevant literature concerns variational inference, and related derivative free sampling approaches.
@@ 1.3.1 Variational Inference
where the minimizer coincides with $\rho_{\rm post}$. The energy function is often chosen as the Kullback-Leibler divergence (4). Along the gradient flow of the energy function, $\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\nabla_{M}\mathcal{E}(\rho,\rho_{\rm post})$, with respect to the metric $M$, the density $\rho$ gradually converges towards the target density. In practice, variational inference methods can be categorized into two categories: non-parametric and parametric approaches.
In non-parametric variational inference, the variational distribution is represented by an interacting particle system. The choice of metrics [13] determines the gradient flow, including the Wasserstein gradient flow [35, 12, 40], Fisher-Rao gradient flow [47, 17, 11, 63], Wasserstein-Fisher-Rao gradient flow [45], Kalman-Wasserstein gradient flow [22] and Stein gradient flow [42]. The interacting particle system evolves according to the gradient flow, gradually approximating the target distribution. The convergence rate and accuracy depend on the number of particles used to represent the distribution and the quality of their representation.
In parametric variational inference, the variational densities used to approximate the target density are parametrized. A common choice for the variational family is the Gaussian, which leads to Gaussian variational inference [50, 40, 5]. To account for the geometric structure of parameters, the gradient is preconditioned by the Fisher information matrix, resulting in the Fisher-Rao natural gradient [1, 50]. This approach leverages Fisher information matrix as a substitute for the Hessian, thereby outperforming standard gradient descent [46]. Gaussian approximations have also been explored for Stein gradient flows [43] and Wasserstein gradient flows [40, 16]. However, the Gaussian approximation has limited expressive power, prompting the use of more flexible variational families, such as Gaussian mixture. These include Gaussian mixture approximation of the natural gradient flow [41, 14] and Wasserstein gradient flow [40, 34]. The present work focuses on Gaussian mixture approximation of the natural gradient flow, with an emphasis on stable and efficient derivative free approximations.
@@ 1.3.2 Derivative Free Sampling Approaches
A large class of derivative free sampling approaches are Markov chain Monte Carlo (MCMC) methods [26, 25], where a derivative free proposal is used to move particles. The main challenges of these methods are the absence of stopping criteria and slow convergence, which worsen as the dimensionality of the problem increases. Several improvements have been proposed to address these issues, such as enhancing the proposal distribution by preserving affine invariance (e.g., the stretch move method [27]) and leveraging parallelization (e.g., with multiple chains [21, 7, 60]). This work also incorporates these two strategies.
Another class of methods is Sequential Monte Carlo (SMC)[18, 56, 4], where particles and their associated weights are updated together using importance sampling. However, resampling is necessary to address issues such as particle degeneracy, sample impoverishment, and instability[19], which arise due to the lack of gradient flow structures.
Finally, in variational inference, a major approach is black-box variational inference [52], which employs stochastic approximation via Monte Carlo methods to estimate gradients. However, the high variance of these gradient estimates often results in unstable updates and slow convergence. To address these challenges, strategies such as incorporating variance reduction techniques and adaptive learning rates to stabilize the gradient estimates have been proposed in [52, 62]. Alternatively, gradient-based variation inference methods, such as Kalman-Wasserstein gradient flow [22] and Fisher-Rao gradient flow [14], avoid direct gradient computation by leveraging Stein’s lemma: $\mathbb{E}_{\mathcal{N}(\theta;m,C)}[\nabla_{\theta}\mathcal{F}(\theta)]=C^{-1}{\rm Cov}_{\mathcal{N}(\theta;m,C)}[\mathcal{F}(\theta),\theta]$. This approach enables gradient estimation through quadrature rules originated from Kalman filtering techniques, such as the ensemble Kalman filter [20], the unscented Kalman filter [37, 36, 33], and the cubature Kalman filter [2]. Notably, the latter two achieve exactness for linear $\mathcal{F}$, indicating low variance. Despite these advancements, these approaches still require small time steps to mitigate instability, especially when the posterior is high-dimensional, exhibits multimodality, or involves complex dependencies. This work focuses on advancing quadrature rules for estimating gradients and even Hessians, which are crucial for developing stable, derivative-free variational inference methods.
@@ 1.4 Organization
The remainder of the paper is organized as follows. In Section 2, we provide an overview of natural gradient variational inference, including both Gaussian and Gaussian mixture variational families. Section 3 introduces our Derivative Free Gaussian Mixture Variational Inference (DF-GMVI), and the related theoretical insights are presented in Section 4. Numerical experiments are described in Section 5, which serve to empirically validate the theory and demonstrate the effectiveness of the proposed framework for Bayesian inference. Finally, concluding remarks are provided in Section 6.
2 Natural Gradient Variational Inference
In this section, we first briefly review natural gradient variational inference from the perspective of gradient flow. Then we discuss both the Gaussian variational family and the Gaussian mixture variational family in Sections 2.1 and 2.2, respectively. Variational inference aims to approximate the posterior distribution Eq. 2 by minimizing the Kullback–Leibler (KL) divergence [61, 5]
@@ 2 Natural Gradient Variational Inference
The steepest descent direction can be interpreted as
[…]
This modification leads to a different gradient flow for updating $a$ as
@@
2.1 Gaussian Approximation
$\rho_{a}^{\rm G}(\theta)=\mathcal{N}(\theta;m,C)$
(11) ${\rm FIM}(a)=\begin{bmatrix}C^{-1}&\\ &X\end{bmatrix},$
(12) $XY=\frac{1}{4}C^{-1}(Y+Y^{T})C^{-1}.$
Incorporating the Fisher information matrix into Eq. 10 and applying Eq. 2 yield the following natural gradient flow
(13) $\dot{m}_{t}=-C_{t}\mathbb{E}_{\rho_{a_{t}}}[\nabla\Phi_{R}],\qquad\dot{C}_{t}=C_{t}-C_{t}\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]C_{t}.$
(14) $\dot{C}_{t}^{-1}$ $=\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]-C_{t}^{-1}.$
The above dynamics Eq. 13 is affine invariant [13, Section 5.4.1]. Consequently, when the posterior is Gaussian, it converges exponentially fast to the posterior at a rate of $\mathcal{O}(e^{-t})$ [13, Theorem 5.6][22, 10, 8], where the exponent of the convergence rate is independent of the posterior. Furthermore, the time discretization of the natural gradient flow Eq. 13 exhibits superior stability. This stability arises from the fact that when the posterior density is log-concave, i.e., when $\nabla^{2}\Phi_{R}$ is positive semidefinite, the forward Euler discretization of Eq. 14 with $10<\Delta t<$ 1 ensures covariance positivity:
$C_{t+\Delta t}^{-1}=C_{t}^{-1}+\Delta t\bigl{(}\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]-C_{t}^{-1}\bigr{)}=(1-\Delta t)C_{t}^{-1}+\Delta t\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]\succ 0,$
provided that the approximation of $\mathbb{E}_{\rho_{a_{t}}}[\nabla^{2}\Phi_{R}]$ remains positive semidefinite.
2.2 Gaussian Mixture Approximation
$\rho_{a}^{\rm GM}(\theta)=\sum_{k=1}^{K}w_{k}\mathcal{N}(\theta;m_{k},C_{k})$
(16) $\operatorname*{arg\,min}_{\sigma}\frac{\langle\nabla_{a}{\rm KL}[\rho_{a}\|\rho_{\rm post}],\sigma\rangle}{\sqrt{\langle\sigma,\textrm{FIM}(a)\sigma\rangle}}\quad\textrm{ s.t. }\quad\sum_{k=1}^{K}\sigma_{\dot{w}_{k}}=0,$
where $\{\sigma_{\dot{w}_{k}}\}$ represents the descent directions corresponding to the weights. The Karush–Kuhn–Tucker conditions of Eq. 16 lead to the following natural gradient flow:
Here $\lambda\in\mathbb{R}$ is the Lagrangian multiplier, determined by the constraint $\sum\dot{w}_{k}=1$. Its value depends on ${\rm FIM}(a)$.
The Fisher information matrix ${\rm FIM}(a)$ for Gaussian mixtures does not have a closed-form expression, and its inversion is computationally expensive. To improve efficiency, diagonal approximations of the Fisher information matrix have been used in the literature [14, Appendix C.8][41], leading to the following approximation:
(19) $X_{k}Y=\frac{1}{4}C_{k}^{-1}(Y+Y^{T})C_{k}^{-1}.$

### v3

@@ Abstract
This paper is concerned with the approximation of probability distributions known up to normalization constants, with a focus on Bayesian inference for large-scale inverse problems in scientific computing. In this context, key challenges include costly repeated evaluations of forward models, multimodality, and inaccessible gradients for the forward model. To address them, we develop a variational inference framework that combines Fisher-Rao natural gradient with specialized quadrature rules to enable derivative free updates of Gaussian mixture variational families. The resulting method, termed Derivative Free Gaussian Mixture Variational Inference (DF-GMVI), guarantees covariance positivity and affine invariance, offering a stable and efficient framework for approximating complex posterior distributions. The effectiveness of DF-GMVI is demonstrated through numerical experiments on challenging scenarios, including distributions with multiple modes, infinitely many modes, and curved modes in spaces with up to 100 dimensions. The method’s practicality is further demonstrated in a large-scale application, where it successfully recovers the initial conditions of the Navier-Stokes equations from solution data at positive times.
@@ 1 Introduction
Sampling a target probability distribution known up to a normalization constant is a classical problem in scientific computing. Specifically, in Bayesian inverse problems [47, 67], the goal is to recover an unknown parameter $\theta\in\mathbb{R}^{N_{\theta}}$ from noisy observation $y\in\mathbb{R}^{N_{y}}$, governed by the equation
@@
Here, $\mathcal{G}$ represents a forward map which, for the problems we consider, is nonlinear and requires solving a partial differential equation (PDE) for each evaluation. The observation noise $\eta$ follows a Gaussian distribution: $\eta\sim\mathcal{N}(0,\Sigma_{\eta})$. Within the Bayesian framework, we assign a Gaussian prior $\mathcal{N}(r_{0},\Sigma_{0})$ to the unknown parameter $\theta$, resulting in a posterior distribution from which we aim to draw samples
@@
We can also express $\Phi_{R}$ through a nonlinear least-squares form, with an augmented map $\mathcal{F}(\theta)$ satisfying
@@ 1.1 Challenges
For many Bayesian inverse problems in scientific applications, computing gradients of $\Phi_{R}$ might be infeasible, as it requires derivatives of large-scale PDE-based models $\mathcal{G}$ that may be black-box (e.g., climate models [65, 76, 32]), use non-differentiable numerical methods (e.g., embedded boundary methods [62, 37, 36, 10] and adaptive mesh refinement [4, 7]), or model discontinuous physics (e.g., in fracture [58] or cloud modeling [68, 53]). Derivative free methods are thus needed; see a review of methodologies in Section 1.3.2.
While there are a few Markov Chain Monte Carlo (MCMC) and Sequential Monte Carlo (SMC) approaches that do not require gradients, they often require numerous function evaluations, particularly in high dimensions, to achieve convergence or mitigate weight collapse. This makes them impractical in many large scale inverse problems given the high computational cost of evaluating $\mathcal{G}$. The computational cost can be reduced by (online) constructing surrogate models [24, 49, 74, 16, 75] and/or employing multifidelity strategies [30, 59, 1] for the forward model $\mathcal{G}$, but additional sources of error might be introduced. Furthermore, the potential multimodality of $\rho_{\rm post}(\theta)$ can cause MCMC methods to struggle with mode transitions [26, 27] and complicates surrogate model training. Missing modes can lead to significant prediction errors in scientific applications [69].
Variational inference offers a promising scalable alternative, with black-box variational inference (BBVI) [63] being a popular derivative free approach. However, BBVI typically relies on stochastic Monte Carlo approximations to estimate gradients, which exhibit high variance. This high variance can make BBVI unreliable, often necessitating variance reduction techniques and extensive time step tuning. Even with these improvements, small time steps remain necessary to maintain stability; see a demonstration in Section 6.2.
On the other hand, Kalman methodology has also been used to develop derivative free variational inference methods [15, 38] based on approximations inspired by ensemble Kalman filters. While these approaches can be effective for posterior distributions close to Gaussian or Gaussian mixtures with well-separated components, they often become inaccurate or even unstable in the presence of non-Gaussian or strongly curved modes; see the experiments reported in [15, Appendix E] and detailed discussion in Section 2.2.
Motivated by these challenges, in this work, we explore guidelines for designing stable derivative free Gaussian mixture variational inference methods. We take the perspective that both BBVI [63] and the Kalman methodology-based approaches [15, 38] are numerically approximating a gradient flow equation. In particular, with natural gradients, the flow equation becomes the Fisher-Rao gradient flow. Our guidelines address stable numerical approximation of the Fisher-Rao gradient flow. Building on the guidelines, we propose a novel approach that requires no hyperparameter tuning while achieving both high accuracy and computational efficiency.
@@ 1.2 Contributions
    We introduce practical guidelines for numerically solving the Fisher-Rao gradient flow within the Gaussian mixture variational family. These guidelines emphasize the consistent integration of entropy and cross-entropy terms to ensure accuracy, and the maintenance of covariance matrix positivity as a critical indicator of numerical stability. They provide a foundation for enhancing the robustness of existing Gaussian mixture variational inference methods and developing new algorithms.
@@
    Building on these guidelines, we propose a derivative free variational inference approach (DF-GMVI), which employs novel quadrature rules to approximate expectations involving Gaussian mixtures and $\Phi_{R}$. These quadrature rules go beyond linear approximations by capturing curvature information of the augmented map $\mathcal{F}$, while maintaining linear computational cost with respect to the dimensionality of $\theta$ and requiring no derivatives. DF-GMVI ensures covariance positivity and exhibits affine invariance, both of which contribute to its superior stability, even with large time steps.
@@
    We demonstrate that DF-GMVI effectively captures multiple or even infinite modes, as well as curved modes, in model problems involving up to 100 dimensions. It also performs well in PDE-based applications, such as reconstructing the Navier-Stokes initial condition from solution data at positive times.
@@ 1.3.1 Variational Inference
where the minimizer coincides with $\rho_{\rm post}$. The energy function is often chosen as the Kullback-Leibler divergence (4). This optimization problem is then addressed via gradient flows and their discretizations. Along the gradient flow of the energy function, $\frac{\mathrm{d}\rho}{\mathrm{d}t}=-\nabla_{M}\mathcal{E}(\rho,\rho_{\rm post})$, with respect to a selected metric $M$, the energy gradually decreases.
In practice, we can classify variational inference methods into two categories: non-parametric and parametric approaches. In non-parametric variational inference, the variational distribution is often represented by a particle system. Gradient flows with various metrics [14] have been used in this context, including the Wasserstein gradient flow [43, 13, 48], Fisher-Rao gradient flow [57, 19, 12, 77], Wasserstein-Fisher-Rao gradient flow [55], Kalman-Wasserstein gradient flow [25] and Stein gradient flow [51, 75, 42]. The particle system evolves according to the gradient flows and gradually approximates the target distribution. The convergence rate and accuracy depend on the number of particles used to represent the distribution and the quality of their representation.
In parametric variational inference, the variational distribution is a parametric density. A common choice is the Gaussian distribution, which leads to Gaussian variational inference [61, 48, 6]. To account for the geometric structure of parameter space, the gradient is preconditioned by the Fisher information matrix, resulting in the Fisher-Rao natural gradient [2, 61] methods. This approach has been shown to outperform standard gradient descent [56]. Researchers have also explored Gaussian approximations for Stein gradient flows [52] and Wasserstein gradient flows [48, 18]. However, the Gaussian approximation offers limited expressive power, which has motivated the use of more flexible variational families, such as Gaussian mixtures. These include Gaussian mixture approximations of the natural gradient flow [50, 15] and Wasserstein gradient flow [48, 40]. Specifically, the present work focuses on Gaussian mixture approximation of the natural gradient flow, emphasizing stable and derivative free approximation methods that enhance computational efficiency.
@@ 1.3.2 Derivative Free Sampling Approaches
A large class of derivative free sampling approaches is based on Markov chain Monte Carlo (MCMC) [29, 28], where a derivative free proposal is used to move particles. The main challenges of these methods are the absence of stopping criteria and slow convergence, which worsen as the dimensionality of the problem increases. Several improvements have been proposed to address these issues, such as enhancing the proposal distribution by preserving affine invariance (e.g., the stretch move method [31]) and leveraging parallelization (e.g., with multiple chains [23, 8, 70]). This work also incorporates these two strategies.
Another class of methods is Sequential Monte Carlo (SMC) [20, 66, 5, 54], where particles and their associated weights are updated together using importance sampling. The importance weights, especially in high dimensions, can suffer from collapse issues , and techniques such as frequent resampling are needed to address particle degeneracy, sample impoverishment, and instability [21].
Finally, in variational inference, a major derivative free approach is black-box variational inference [63], which employs stochastic approximation via Monte Carlo methods to estimate gradients. However, the high variance of these gradient estimates often results in unstable updates and slow convergence. To address these challenges, researchers have proposed strategies such as incorporating variance reduction techniques and adaptive learning rates to stabilize gradient estimates [63, 72]. Alternatively, gradient-based variational inference methods, such as Kalman-Wasserstein gradient flow [25] and Stein gradient flow [34], avoid direct gradient computation by leveraging Stein’s lemma:
$\mathbb{E}_{\mathcal{N}(\theta;m,C)}[\nabla_{\theta}\mathcal{F}(\theta)]=C^{-1}{\rm Cov}_{\mathcal{N}(\theta;m,C)}[\mathcal{F}(\theta),\theta].$
This approach enables gradient estimation through quadrature rules derived from Kalman filtering techniques, such as the ensemble Kalman filter [22], the unscented Kalman filter [45, 44, 39], and the cubature Kalman filter [3]. Notably, the latter two achieve exactness for linear $\mathcal{F}$. Despite these advancements, these approaches still require small time steps to mitigate instability, especially when the posterior is high-dimensional, exhibits multimodality, or involves complex dependencies. The present work advances quadrature rules for estimating gradients and Hessians, which are crucial for developing stable, derivative free variational inference methods.
@@ 1.4 Organization
In Section 2, we provide an overview of natural gradient variational inference, with a focus on Gaussian mixture variational families. Section 3 outlines practical guidelines for numerically solving natural gradient flow. Section 4 introduces our Derivative Free Gaussian Mixture Variational Inference (DF-GMVI), and the related theoretical insights are presented in Section 5. Numerical experiments are described in Section 6, which serve to empirically validate the theory and demonstrate the effectiveness of the proposed framework for Bayesian inference. Finally, concluding remarks are provided in Section 7.
2 Variational Inference with Natural Gradient
In this section, we briefly review variational inference with natural gradient methods, from the perspective of gradient flows. We focus on Gaussian mixture variational families.
In detail, we approximate the posterior distribution Eq. 2 by minimizing the Kullback–Leibler (KL) divergence [71, 6]
@@ 2 Variational Inference with Natural Gradient
over a family of variational densities $\rho_{a}$, parameterized by $a\in\mathbb{R}^{N_{a}}$. In Gaussian mixture variational inference, the variational family is a $K$-component Gaussian mixture
$\rho_{a}^{\rm GM}(\theta)=\sum_{k=1}^{K}w_{k}\mathcal{N}(\theta;m_{k},C_{k}),$
parameterized by means $m_{k}\in\mathbb{R}^{N_{\theta}}$, covariances $C_{k}\in\mathbb{R}^{N_{\theta}\times N_{\theta}}$ and weights $w_{k}\in\mathbb{R}_{\geq 0}$, collectively denoted by the parameter vector
$a:=[m_{1},...,m_{k},...,m_{K},C_{1},...,C_{k},...,C_{K},w_{1},...,w_{k},...,w_{K}].$
Weights satisfy $\sum_{k=1}^{K}w_{k}=1$.
The minimization of (4) can be done by using gradient flows. The natural gradient method [2] employs the Fisher information matrix [64] as a preconditioner:
This leads to the following gradient flow for updating $a$:
@@
This equation can also be understood as a Gaussian mixture approximation of the Fisher-Rao gradient flow in the space of probability densities [14].
2.1 Approximate Fisher Information
For the Gaussian mixture variational family, the Fisher information matrix lacks a close-form expression, and its inversion is computationally challenging. To address this, following [15, Appendix C.8][50], we approximate the Fisher information matrix with a block-diagonal form. Substituting this approximation into Eq. 5 leads to the following equations: